# 1]Independent Researcher

### v1

@@ 3 Results and discussions
In this section, we apply the proposed PINN framework to the (2+ 1)-dimensional VC-CNLSE and assess its accuracy and performance in predicting various soliton dynamics. First, we address the problem of data-driven soliton solutions to the model equation, predicting phenomena such as vector parabolic dark one-soliton, vector m-shaped anti-dark one-soliton, and dark two-solitons. By considering different formations of dispersion and nonlinearity coefficients, we demonstrate that our method is capable of predicting the corresponding soliton dynamics with high accuracy. Second, we explore the parameter discovery within the model equation, focusing on accurately identifying both constant and variable coefficients.
@@ 3.1 Data-driven solution to (2+ 1)-dimensional VC-CNLSE
Choosing different and appropriate formations of dispersion and nonlinearity coefficients leads us to explore a wide range of soliton dynamics and behaviors. By adjusting these coefficients, we can model various physical scenarios and validate the robustness and versatility of our proposed PINN framework. This capability highlights the framework’s potential to be applied to complex, real-world problems where the underlying physics is governed by variable coefficients. In doing so, we consider the equation system (1) equipped with the initial condition (5) and the boundary condition (6).
3.1.1 Discovery of one-soliton solutions
@@
$\beta(t)$ $:=-\frac{\csc^{2}(\theta)(k^{2}+m^{2})\alpha(t)}{2(\lvert\lambda\rvert^{2}+\lvert\mu\rvert^{2})},$ (14)
[…]
$\xi(x,y,t)$ $:=kx+my+w(t)+\delta.$ (16)
@@
Here, $k$, $m$, and $\delta$ are real constants. $\lambda$ and $\mu$ are complex constants, while $\theta$ is a real phase shift. One can see that the one-soliton solutions (12) are proportional to each another. This leads us to dealing with a Manakov system when we set $\tau=1$ in the system of equations (1).
  • Parabolic vector dark one-soliton
    Based on the solutions (12), the parabolic vector dark one-soliton for the (2+ 1)-dimensional VC-CNLSE (1) can be obtained by taking $k=-1.4,m=1.1,\delta=1,\mu=-2,\theta=3.8,\lambda=-2$, and $\alpha(t)=t$. With this configuration, one can see that the amplitudes for both $\psi_{1}$ and $\psi_{2}$ are the same. We set the spatial domain $\Omega=[-10,10]\times[-10,10]$ and the temporal domain $[-1,1]$. Then, the initial and boundary conditions on these regions can be obtained as:
    $\begin{cases}\psi_{j}(x,y,t=-1)=-2\frac{e^{ib(-1)}(1+e^{\xi(x,y,-1)+i7.6})}{1+e^{\xi(x,y,-1)}},\quad(x,y)\in\Omega,\\ \psi_{j}(x=-10,y,t)=-2\frac{e^{ib(t)}(1+e^{\xi(-10,y,t)+i7.6})}{1+e^{\xi(-10,y,t)}},\quad(y,t)\in[-10,10]\times(-1,1],\\ \psi_{j}(x,y=-10,t)=-2\frac{e^{ib(t)}(1+e^{\xi(x,-10,t)+i7.6})}{1+e^{\xi(x,-10,t)}},\quad(x,t)\in[-10,10]\times(-1,1],\\ \psi_{j}(x=10,y,t)=-2\frac{e^{ib(t)}(1+e^{\xi(10,y,t)+i7.6})}{1+e^{\xi(10,y,t)}},\quad(y,t)\in[-10,10]\times(-1,1],\\ \psi_{j}(x,y=10,t)=-2\frac{e^{ib(t)}(1+e^{\xi(x,10,t)+i7.6})}{1+e^{\xi(x,10,t)}},\quad(x,t)\in[-10,10]\times(-1,1],\quad j=1,2.\\ \end{cases}$
    Based on the information above, we first generate the training dataset corresponding to the initial and boundary conditions. We randomly generate $N_{ic}=100$ initial points and $N_{bc}=100$ boundary points. Beyond that, we divide the computational domain $\Omega\times[-1,1]$ into $256\times 256\times 201$ discrete equidistant points to test the accuracy of the underlying PINN model. In order to inform the physics models (10) into the neural network’s knowledge, we use the LHS technique to generate $N_{r}=10,000$ residual points. These points are used to penalize the physics models at the so-called collocation points. Then, the overall loss function can be constructed by considering the MSE losses corresponding to the initial and boundary conditions, the physics models, and the loss associated with the slope recovery. For this problem, we consider a neural network model with $5$ hidden layers and $40$ neurons in each layer. We start by training our model with $30,000$ iterations of the Adam optimizer with the default learning rate to minimize the corresponding loss function. Then, we continue to obtain the optimal values of the network’s parameters by performing $5,144$ iterations of the L-BFGS optimizer. The training time was recorded as $1729.6354\,s$ and $780.5243\,s$ for the Adam optimizer and the L-BFGS optimizer, respectively. After $30,000$ iterations of the Adam optimizar, the $\mathbb{L}^{2}$ relative errors for the solutions $\psi_{1}$ and $\psi_{2}$ are measured as $0.18\%$ and $0.19\%$. By continuing the optimization using the L-BFGS optimizer, the $\mathbb{L}^{2}$ relative errors decrease to $0.073\%$ and $0.064\%$, respectively. After all, we are able to plot the obtained results from the trained PINN model.
@@
    Using the RAR strategy, we first train the underlying model with $N_{f}=10,000$ residual points and then gradually add $6,200$ new points during the training. As depicted in the first column of Fig. 2, the added residual points corresponding to the residual equations have different distributions. It means that each residual equation in (10) has its own characteristics, and this importance should be considered when applying PINN methods to such model equations. The three-dimensional plots of point-wise errors for the solution $\psi_{1}$ along the $x-y$ axis at two temporal levels, $t=-1$ and $t=1$, are depicted in the second and third columns of Fig. 2, respectively. One can see that using RAR drastically decreases the prediction errors and leads residual equations to have uniform distributions of residual points. This phenomenon can be addressed by forcing the residuals to zero during training. In Fig. 3, the density plots of the exact and predicted dynamics, as well as the comparison between the exact and predicted solutions for $\psi_{1}$ at different temporal levels, are depicted. In the top row, the density plots along the $x-t$ plane at $y=0$ and $y-t$ plane at $x=0$ are shown, where the pulse propagates parabolically with opposite directions in $x-t$ and $y-t$ planes. In the middle row, the density plots of the exact and predicted dynamics at two temporal levels, $t=-1$ and $t=1$, along the $x-y$ plane are depicted. One can observe that the pulse propagates along a straight line. A comparison between the exact and predicted solutions at four temporal levels for $x=10$ and $y=10$ is presented in the bottom row. These results show that the trained PINN model can accurately capture the exact dynamics in the $x-t$, $y-t$, and $x-y$ planes, demonstrating the robustness of our method. The evolution of the parabolic dark one-soliton in the $x-t$ plane (first row) at $y=-8$, $y=0$, and $y=8$ (from left to right) and in the $y-t$ plane (bottom row) at $x=-8$, $x=0$, and $x=8$ (from left to right) is illustrated in Fig. 4. The behavior of the loss functions during training with $30,000$ iterations of the Adam optimizer and $5,144$ iterations of the L-BFGS optimizer is depicted in Fig. 5. From the left panel one can see that the Adam optimizer exhibits a slowly decreasing loss with several oscillations during training. The loss values after training with the Adam optimizer are recorded as follows: $21.337e-$ 2 for the overall loss function, $41.732e-$ 4 for the initial condition loss function, $43.283e-$ 4 for the boundary condition loss function, and $36.014e-$ 3 for the residual loss function. From the middle panel of Fig. 5 one can see that the oscillatory behavior of the MSE loss for each individual residual equation is relatively close to others. In contrast, the L-BFGS optimizer exhibits smoother behavior of the loss curves without any oscillations as depicted in the right panel of Fig. 5. The loss values after training with the L-BFGS optimizer are recorded as follows: $038.483e-$ 03 for the overall loss function, $64.587e-$ 6 for the initial condition loss function, $66.705e-$ 6 for the boundary condition loss function, and $59.559e-$ 5 for the residual loss function. It should be noted that, in comparison with the results obtained from the Adam optimizer, the L-BFGS optimizer demonstrates a faster convergence framework.
@@
    Figure 4: Three-dimensional plots with contour map of the predicted dynamics for $\psi_{j}$ in the $x-t$ (top row) and $y-t$ (bottom row) planes.
@@
  • Vector m-shaped anti-dark one-soliton
    Taking $k=-0.19$, $m=-1.5$, $\delta=1$, $\mu=-1.1+i$, $\theta=-4.1$, $\lambda=-1.1+i$, and $\alpha(t)=t\ln t^{2}$ in (12)-(16), one can obtain the vector m-shaped anti-dark one-soliton solution to the (2+ 1)-dimensional VC-CNLSE [25]. Based on these configurations, one can see that the amplitudes for both $\psi_{1}$ and $\psi_{2}$ are the same. We set the spatial domain as $\Omega=[-25,25]\times[-4,4]$ and the temporal domain as $[-2,2]$. In order to generate the training data corresponding to the initial and boundary conditions, for $j=1,2$ we consider the following equations:
    $\begin{cases}\psi_{j}(x,y,t=-2)=(-1.1+i)\frac{e^{ib(-2)}(1+e^{\xi(x,y,-2)-i8.2})}{1+e^{\xi(x,y,-2)}},\quad(x,y)\in\Omega,\\ \psi_{j}(x=-25,y,t)=(-1.1+i)\frac{e^{ib(t)}(1+e^{\xi(-25,y,t)-i8.2})}{1+e^{\xi(-25,y,t)}},\quad(y,t)\in[-4,4]\times(-2,2],\\ \psi_{j}(x,y=-4,t)=(-1.1+i)\frac{e^{ib(t)}(1+e^{\xi(x,-4,t)-i8.2})}{1+e^{\xi(x,-4,t)}},\quad(x,t)\in[-25,25]\times(-2,2],\\ \psi_{j}(x=25,y,t)=(-1.1+i)\frac{e^{ib(t)}(1+e^{\xi(25,y,t)-i8.2})}{1+e^{\xi(25,y,t)}},\quad(y,t)\in[-4,4]\times(-2,2],\\ \psi_{j}(x,y=4,t)=(-1.1+i)\frac{e^{ib(t)}(1+e^{\xi(x,4,t)-i8.2})}{1+e^{\xi(x,4,t)}},\quad(x,t)\in[-25,25]\times(-2,2].\\ \end{cases}$
@@
    We divide the computational domain $\Omega\times[-2,2]$ into $256\times 256\times 200$ discrete equidistant points to create a dataset for testing the performance of the underlying PINN model. We randomly generate $N_{ic}=100$ and $N_{bc}=100$ training points corresponding to the initial and boundary conditions, respectively. We also use the LHS technique to generate $N_{r}=10,000$ residual points within the computational domain. Then, a PINN model based on the enhanced framework, equipped with a neural network consisting of $5$ hidden layers and $40$ neurons per layer, is going to be trained with $N_{r}=10,000$ residual points. Using the RAR strategy, we gradually add $6,200$ residual points during training. After $30,000$ iterations with the Adam optimizer and $5,170$ iterations with the L-BFGS optimizer, we are ready to recover the m-shaped anti-dark one-soliton solutions using the trained model. The $\mathbb{L}^{2}$ relative errors for the solutions $\psi_{1}$ and $\psi_{2}$ are reduced from $6.4\%$ and $6.4\%$ to $0.29\%$ and $0.28\%$, respectively, by using the Adam optimizer and then continuing the optimization process with the L-BFGS optimizer. The training times are recorded as $1412.4702\,s$ and $740.1036\,s$ for the Adam and L-BFGS optimizers, respectively. In Fig. 6, the density plots of the exact and predicted soliton dynamics are shown in the top and middle rows. It can be seen from the top row that our model is able to successfully capture the m-shaped dynamics of the anti-dark one-soliton solution. These dynamics exhibit an m-shaped pulse propagation with two smooth crests in the $x-t$ and $y-t$ planes. One can observe that the two crests in the $x-t$ plane are wider than those in the $y-t$ plane. On the other hand, the pulse propagates in the $x-y$ plane along a straight line, as depicted in the middle row. A comparison between the exact and predicted dynamics is presented at different temporal levels for $y=4$ and $x=25$ in the bottom row to show the accuracy of the trained model. The evolution of the m-shaped anti-dark one-soliton solution and the corresponding point-wise errors in the $x-t$ and $y-t$ planes are depicted in Fig. 7 for $x=0$ and $y=0$ through three-dimensional plots. In Fig. 8, the loss curves during training with $30,000$ iterations of the Adam optimizer and $5,170$ iterations of the L-BFGS optimizer are illustrated. As expected, the Adam optimizer exhibits a slow decrease in loss values with many oscillations during training. From the middle panel of Fig. 8, the MSE loss for each individual residual equation oscillates relatively close to one another. On the other hand, the loss values exhibit smooth behavior without any oscillations when using the L-BFGS optimizer, as shown in the right panel of Fig. 8. The loss values after training with the Adam optimizer are recorded as follows: $021.521e-$ 02 for the overall loss function, $031.118e-$ 03 for the initial condition loss function, $043.675e-$ 04 for the boundary condition loss function, and $035.034e-$ 03 for the residual loss function. The loss values after training with the L-BFGS optimizer are recorded as follows: $038.881e-$ 03 for the overall loss function, $062.622e-$ 06 for the initial condition loss function, $051.184e-$ 05 for the boundary condition loss function, and $042.571e-$ 04 for the residual loss function. These results show that the L-BFGS optimizer achieves lower loss values across all categories compared to the Adam optimizer. Specifically, the L-BFGS optimizer demonstrates superior performance in minimizing the residual loss function and initial condition loss function, reflecting its efficiency in converging to a more accurate solution with fewer oscillations. This indicates that the L-BFGS optimizer provides a more stable and effective framework for training the PINN model in this context.
@@
    Figure 7: Three-dimensional plots with contour maps and the corresponding point-wise errors for $\psi_{j}$ in the $x-t$ plane (first and second panels) and $y-t$ plane (third and fourth panels), from left to right.

### v2

@@ 3 Results and discussions
In this section, we apply the proposed PINN framework to the (2+ 1)-dimensional VC-CNLSE and assess its accuracy and performance in predicting various soliton dynamics. First, we address the problem of data-driven soliton solutions to the model equation, predicting phenomena such as vector parabolic dark one-soliton, vector m-shaped anti-dark one-soliton, and dark two-solitons. By considering different formations of dispersion and nonlinearity coefficients, we demonstrate that our method is capable of predicting the corresponding soliton dynamics with high accuracy. Second, we explore the parameter discovery within the model equation, focusing on accurately identifying both constant and variable coefficients. In all our experiments, we initialize the scalable parameters such that $n=5$ and $a_{k}^{\ell}=0.2$ for $\ell=1,\cdots,D-1$ and $k=1,\cdots,N^{\ell}$ and initialize the network’s parameters using the Xavier initialization technique. We set the initial condition loss, boundary condition loss, and slope recovery loss weights as: $w_{ic}=1$, $w_{bc}=1$, and $w_{a}=\frac{1}{100}$. Setting $w_{a}=\frac{1}{100}$ ensures that the value of the loss is not too large. We generate the residual points randomly using the Latin Hypercube Sampling (LHS) strategy [55] within the computational domain. We start training the constructed PINN model with an initial number of residual points, setting $m=25$. After every $5000$ iterations, we add $2^{k}\times m$ new residual points for each residual equation based on the RAR algorithm during the training phase, where $k$ denotes the number of times the $5000$-iteration update has occurred, and this sampling process is continued until either the maximum number of iterations is reached or the mean equation residuals fall below 1e61e-6. We set the stopping criterion to $\delta=1e-05$ during the training of the underlying PINN model. The accompanying code for this paper is written in Python, and the numerical results are reported using the JAX library. The JAX library is used for its efficient automatic differentiation and high-performance numerical computing capabilities. The computations were performed on a Lenovo Legion laptop equipped with a 2.30 GHz Intel Core i7-11800H processor and a single NVIDIA GeForce RTX 3060 GPU card. To evaluate the performance and accuracy of our method, we consider two metrics, the $\mathbb{L}^{2}$ error and the point-wise error, defined as follows:
where $\psi_{j}(\cdot,\cdot)$ and $\hat{\psi}_{j,\boldsymbol{\theta}}(\cdot,\cdot)$ are the exact and predicted solutions, respectively, evaluated at sampled test points.
@@ 3.1 Data-driven solution to (2+ 1)-dimensional VC-CNLSE
Choosing different and appropriate formations of dispersion and nonlinearity coefficients leads us to explore a wide range of solitonic dynamics. By adjusting these coefficients, we can model various physical scenarios. Here, we aim to employ the proposed PINNs framework based on Algorithm 1, considering the equation system (1) equipped with the initial condition (3) and the boundary condition (4) for data-driven solution of the (2+ 1)-dimensional VC-CNLSE. For each experiment, the network configurations for the underlying PINN model and the hyper-parameter settings are reported in Table 6.
Table 1: Network configurations and hyper-parameter settings for experimental cases.
Network
Cases $N_{ic}$ $N_{bc}$ Iterations Learning rate Depth Width
1 $100$ $100$ $35,144$ $0.001$ 55 $40$
2 $100$ $100$ $35,170$ $0.001$ 55 $40$
3 $200$ $200$ $35,215$ $0.0001$ 44 $100$
4 $200$ $200$ $35,034$ $0.0001$ 44 $100$
3.2 Discovery of one-soliton solutions
@@
$b(t)$ $:=(\lvert\lambda\rvert^{2}+\lvert\mu\rvert^{2})\int\beta(t)dt,\quad\beta(t):=-\frac{\csc^{2}(\theta)(k^{2}+m^{2})\alpha(t)}{2(\lvert\lambda\rvert^{2}+\lvert\mu\rvert^{2})},$ (11)
$w(t)$ $:=\cot(\theta)(k^{2}+m^{2})\int\alpha(t)dt,\quad\xi(x,y,t):=kx+my+w(t)+\epsilon.$ (12)
@@
Here, $k$, $m$, and $\epsilon$ are real constants. $\lambda$ and $\mu$ are complex constants, while $\theta$ is a real phase shift. One can see that the one-soliton solutions (10) are proportional to each other. This leads us to dealing with a Manakov system when we set $\tau=1$ in the system of equations (1).
Case 1: Parabolic vector dark one-soliton
@@
Based on solutions (10), a parabolic vector dark one-soliton for the (2+ 1)-dimensional VC-CNLSE (1) is obtained with parameters k=1.4k=-1.4, $m=1.1$, $\epsilon=1$, $\mu=-2$, $\theta=3.8$, $\lambda=-2$, and $\alpha(t)=t$, yielding equal amplitudes for $\psi_{1}$ and $\psi_{2}$. We set the spatial domain $\Omega=[-10,10]\times[-10,10]$ and the temporal domain [1,1][-1,1]. We divide the computational domain $\Omega\times[-1,1]$ into $256\times 256\times 201$ discrete equidistant points to randomly generate $N_{ic}=N_{bc}=100$ initial and boundary training points based on the information above. We use the LHS technique to generate $N_{r}=10,000$ residual points in the first stage of training and then gradually add new residual points based on Algorithm 1. For this problem, we consider a neural network model with $5$ hidden layers and $40$ neurons in each layer. The network parameters are optimized using $30,000$ iterations of the Adam optimizer followed by $5,144$ iterations of the L-BFGS optimizer, reducing the $\mathbb{L}^{2}$ relative errors for $\psi_{1}$ and $\psi_{2}$ from $0.18\%$ and $0.19\%$ to $0.073\%$ and $0.064\%$, respectively. As depicted in the first column of Fig. 2, the added residual points corresponding to the residual equations have different distributions, demonstrating that each residual equation exhibits distinct characteristics, highlighting the need to consider their individual importance when applying PINN methods. The three-dimensional plots of point-wise errors for the solution $\psi_{1}$ along the xyx-y axis at two temporal levels, t=1t=-1 and $t=1$, are depicted in the second and third columns of Fig. 2, respectively. One can see that incorporating RAR significantly reduces prediction errors and yields a more uniform distribution of residual points by driving the residuals towards zero during training. In Fig. 3, the density plots of the exact and predicted dynamics, as well as the comparison between the exact and predicted solutions for $\psi_{1}$ at different temporal levels, are depicted. In the top row, the density plots along the xtx-t plane at $y=0$ and the yty-t plane at $x=0$ are shown, where the pulse propagates parabolically in opposite directions in the xtx-t and yty-t planes. In the middle row, the density plots of the exact and predicted dynamics at two temporal levels, t=1t=-1 and $t=1$, along the xyx-y plane are depicted. One can observe that the pulse propagates along a straight line. A comparison between the exact and predicted solutions at four temporal levels for $x=10$ and $y=10$ is presented in the bottom row. These results show that the trained PINN model can accurately capture the exact dynamics in the xtx-t, yty-t, and xyx-y planes, demonstrating the robustness of our method. The evolution of the parabolic dark one-soliton in the xtx-t plane (first row) at y=8y=-8, $y=0$, and $y=8$ (from left to right) and in the yty-t plane (bottom row) at x=8x=-8, $x=0$, and $x=8$ (from left to right) is illustrated in Fig. 4. The loss trajectories associated with the Adam optimizer and the L-BFGS optimizer are depicted in Fig. 5. From the left panel, one can see that the Adam optimizer exhibits a slowly decreasing loss with several oscillations during training. From the middle panel of Fig. 5, one can see that the oscillatory behavior of the MSE loss for each individual residual equation is relatively close to that of the others. In contrast, the L-BFGS optimizer exhibits smoother behavior of the loss curves without any oscillations as depicted in the right panel of Fig. 5. It should be noted that, in comparison with the results obtained from the Adam optimizer, the L-BFGS optimizer demonstrates a faster convergence framework. Table 2 compares the performance of the proposed enhanced PINN method, the adaptive activation function-based PINN method in [50], and the standard PINN method in terms of $\mathbb{L}^{2}$ relative error, loss values (Adam followed by L-BFGS), and training time (in seconds). For a fair comparison, all networks are configured with $5$ hidden layers and $40$ neurons per layer. For the method in [50], the adaptive activation functions are set the same as in the enhanced PINN. The number of residual points is set to $N_{r}=16,200$ for both the method in [50] and the standard PINN. These results show that the proposed enhanced PINN method outperforms both the PINN method in [50] and the standard PINN method by achieving the lowest $\mathbb{L}^{2}$ relative errors and loss values, along with a shorter training time. This improvement is due to its region-specific training approach, which focuses on areas with high equation residuals and allocates different distributions of residual points for each residual equation, resulting in more effective penalization of the equation residuals during training while using a smaller number of residual points. In contrast, the other methods use large datasets with the same distribution for all residual equations, leading to higher computational cost and longer training times due to the use of the AD technique, which requires computing the derivatives of the network outputs with respect to input data at all residual points during training. Overall, the enhanced PINN demonstrates superior accuracy and computational efficiency for solving the (2+ 1)-dimensional VC-CNLSE problem.
@@
Figure 4: Three-dimensional plots with contour map of the predicted dynamics for $\psi_{j}$ in the xtx-t (top row) and yty-t (bottom row) planes.
@@
Table 2: Comparison results in terms of the $\mathbb{L}^{2}$ relative error, loss value, and training time (s).
$\mathbb{L}^{2}$ error Loss value
Method $\psi_{1}$ $\psi_{2}$ Adam L-BFGS Training time (s)
Enhanced PINN $073\% 0.0$$064\%$ 1.3371e021.3371e-02 8.4830e038.4830e-03 $2510.1597$
Method in [50] $0.37\%$ $0.55\%$ 9.4780e019.4780e-01 8.1209e028.1209e-02 $2741.0294$
PINN $5.12\%$ $4.28\%$ 3.0813e003.0813e-$0$ 6.3085e016.3085e-01 $2571.8029$
Case 2: Vector m-shaped anti-dark one-soliton
@@
By setting the parameters as k=0.19k=-0.19, m=1.5m=-1.5, $\epsilon=1$, $\mu=-1.1+i$, $\theta=-4.1$, $\lambda=-1.1+i$, and $\alpha(t)=t\ln t^{2}$ in equations (10)-(12), a vector m-shaped anti-dark one-soliton solution to the (2+ 1)-dimensional VC-CNLSE is obtained, where both $\psi_{1}$ and $\psi_{2}$ have the same amplitudes. The spatial domain is $\Omega=[-25,25]\times[-4,4]$ and the temporal domain is [2,2][-2,2], discretized into $256\times 256\times 200$ equidistant points. For training, $N_{ic}=100$ initial condition points and $N_{bc}=100$ boundary condition points are randomly generated, while $N_{r}=10,000$ residual points are sampled using the LHS technique. An enhanced PINN model with $5$ hidden layers and $40$ neurons per layer is then trained using these $10,000$ residual points. The $\mathbb{L}^{2}$ relative errors for the solutions $\psi_{1}$ and $\psi_{2}$ are reduced from $6.4\%$ and $6.4\%$ to $0.29\%$ and $0.28\%$, respectively, by using the Adam optimizer and then continuing the optimization process with the L-BFGS optimizer. In Fig. 6, the density plots of the exact and predicted soliton dynamics are shown in the top and middle rows, respectively. The top row demonstrates that the model successfully captures the m-shaped dynamics of the anti-dark one-soliton solution, which exhibits an m-shaped pulse with two smooth crests in both the xtx-t and yty-t planes. Notably, the crests in the xtx-t plane are wider than those in the yty-t plane. In the xyx-y plane, shown in the middle row, the pulse propagates along a straight line. The bottom row presents a comparison between the exact and predicted dynamics at different temporal levels for $y=4$ and $x=25$, highlighting the accuracy of the trained model. The evolution of the m-shaped anti-dark one-soliton solution and the corresponding point-wise errors in the xtx-t and yty-t planes are depicted in Fig. 7 for $x=0$ and $y=0$ through three-dimensional plots. In Fig. 8, the loss curves during training with $30,000$ iterations of the Adam optimizer and $5,170$ iterations of the L-BFGS optimizer are illustrated. As expected, the Adam optimizer exhibits a slow decrease in loss values with many oscillations during training. From the middle panel of Fig. 8, the MSE loss for each individual residual equation oscillates relatively close to one another. On the other hand, the loss values exhibit smooth behavior without any oscillations when using the L-BFGS optimizer, as shown in the right panel of Fig. 8. This indicates that the L-BFGS optimizer provides a more stable and effective framework for training the PINN model in this context. Table 3 shows that the proposed enhanced PINN method achieves the lowest $\mathbb{L}^{2}$ relative errors of $0.29\%$ for $\psi_{1}$ and $0.28\%$ for $\psi_{2}$, outperforming both the method in [50] and the standard PINN. All networks have the same structure and we set $N_{r}=16,200$ for the method in [50] and the standard PINN. The enhanced PINN attains the lowest loss values by optimizing the network parameters with the Adam optimizer, followed by the L-BFGS optimizer. Furthermore, its training time is shorter than that of the method in [50] and similar to the standard PINN despite its much higher accuracy. This efficiency is due to its region-specific training approach with fewer residual points and different distributions per equation residual, whereas the other methods use larger residual point datasets with the same distribution, increasing computational cost. Fig. 9 illustrates the evolution of the $\mathbb{L}^{2}$ relative error of $\psi_{1}$ for training the enhanced PINN model with varying network depth (left panel) and width (right panel). As shown in the left panel, increasing the network depth initially reduces the $\mathbb{L}^{2}$ relative error, but further increases lead to slight error growth, suggesting overfitting or optimization challenges with very deep architectures. In contrast, increasing the network width (right panel) consistently decreases the error up to a certain width, after which it stabilizes. Fig. 10 shows the effect of the number of residual points $N_{r}$, initial condition points $N_{ic}$, and boundary condition points $N_{bc}$ on the $\mathbb{L}^{2}$ relative error of $\psi_{1}$ for training the enhanced PINN model. Increasing $N_{r}$ (left panel) significantly reduces the error, indicating its strong influence on enforcing the PDE. Increasing $N_{ic}$ (middle panel) also decreases the error, improving initial condition enforcement. In contrast, increasing $N_{bc}$ (right panel) has a smaller impact on error reduction. Overall, the number of residual points has the greatest effect on model accuracy.
@@
Figure 7: Three-dimensional plots with contour maps and the corresponding point-wise errors for $\psi_{j}$ in the xtx-t plane (first and second panels) and yty-t plane (third and fourth panels), from left to right.
@@
Table 3: Comparison results in terms of the $\mathbb{L}^{2}$ relative error, loss value, and training time (s).
$\mathbb{L}^{2}$ error Loss value
Method $\psi_{1}$ $\psi_{2}$ Adam L-BFGS Training time (s)
Enhanced PINN $0.29\%$ $0.28\%$ 1.5215e021.5215e-02 8.8810e038.8810e-03 $2152.5738$
Method in [50] $8.03\%$ $8.05\%$ 2.0385e002.0385e-$0$ 8.8399e018.8399e-01 $2403.9105$
PINN $15.19\%$ $15.18\%$ $9.8001e+01$ 7.0091e007.0091e-$0$ $2152.5801$
Figure 9: The effect of the depth and width of the network on the $\mathbb{L}^{2}$ relative error of $\psi_{1}$.
Figure 10: The effect of the number of training and residual points on the $\mathbb{L}^{2}$ relative error of $\psi_{1}$.